OpenStudy (anonymous):

Eliminate the parameter. x = square root of t, y = 3t + 7 y = 3x2 + 7 y = 3square root of x + 7, x ≥ 0 y = 3square root of x - 7, x ≥ 0 y = 3x2 - 7

OpenStudy (anonymous):

It should be y= 3x^2+7

OpenStudy (anonymous):

can u help me with another one? @dpasingh

OpenStudy (anonymous):

Go ahead

OpenStudy (anonymous):

Eliminate the parameter. x = 4 cos t, y = 4 sin t

OpenStudy (anonymous):

Ok pls wait

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x = 4 cos t, y = 4 sin t Squaring each relation separately and then adding them, we find the following \[i.e. x ^{2} = (4 \cos t)^{2}, y^{2} = (4 \sin t)^{2}\] \[x ^{2} + y^{2} = 16 \cos^{2} t+ 16 \sin ^{2}t= 16 (\cos^{2} t+ \sin ^{2}t)=16\times1=16\] \[i.e. x ^{2} + y^{2} = 16\]

OpenStudy (anonymous):

thank you so much!! :)